OpenStudy (anonymous):

the maximum number of volts,E, that can be placed across a resistor is given by the formula E=sq root PR, where P is the number of watts of power that the resistor can absorb and R is the resistance of the resistor in ohms. find Eif P=1/8 watts and R= 12800 ohms A20 volts B. 40 volts C. 1600 volts D. 80 volts

OpenStudy (campbell_st):

just substitute your given values \[E = \sqrt{\frac{1}{8} \times 2800}\] and evaluate

OpenStudy (anonymous):

E = sqrt(PR) which is\[E = \sqrt{PR}\]Just multiply P by R first, then take square root.

OpenStudy (anonymous):

"1" in front of the "2800"

OpenStudy (anonymous):

i have tried everythign nothign is coming close to the answer

OpenStudy (campbell_st):

sorry for the typo try \[E = \sqrt{\frac{1}{8} \times 12800}\]

OpenStudy (campbell_st):

the answer is in the list.

OpenStudy (anonymous):

ok trying again

OpenStudy (anonymous):

the answer 40

OpenStudy (campbell_st):

thats what i got

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

there were 3 problems out of 75 i could not find two left

OpenStudy (campbell_st):

good luck with them